### v2

# An Alignment Between the CRA’s Essential Requirements and the ATT&CK®’s Mitigations

Abstract
The paper presents an alignment evaluation between the mitigations present in the MITRE’s ATT&CK® framework and the essential cyber security requirements of the recently introduced Cyber Resilience Act (CRA) in the European Union. In overall, the two align well with each other. With respect to the CRA, there are notable gaps only in terms of data minimization, data erasure, and vulnerability coordination. In terms of the ATT&CK® framework, gaps are present only in terms of threat intelligence, training, out-of-band communication channels, and residual risks. The evaluation presented contributes to narrowing of a common disparity between law and technical frameworks.
I Introduction
The CRA [1] was agreed upon in 2024. The enforcement will start in 2027. The CRA is a new product-specific cyber security law covering most products with a network connection functionality; the notable exclusions are medical devices, motor vehicles, ships and maritime equipment, and cloud computing. Although stricter obligations are imposed upon products categorized as important or critical, all products must comply with the CRA’s essential cyber security requirements.
To help in bridging a gap between legal and technical domains, the paper investigates an alignment between the CRA’s essential cyber security requirements and the ATT&CK® framework [2]. It is a comprehensive empirical catalog of real-world adversarial tactics and techniques maintained by the non-profit MITRE corporation. It was initiated in 2013. Although the framework’s focus is on advanced persistent threats, the mitigations offered in the framework generalize to countering other threat actors as well. Perhaps partially due to this generalizability, the ATT&CK® framework has been extensively used also in academic research [3, 4, 5]. Among this research is also a previous work for using the framework to elicit security requirements [6]. Another related work worth explicitly mentioning is about the ATT&CK®’s alignment with standards [7]. The paper aligns with and contributes to this line of elicitation and evaluation research.
Recently, the alignment of the CRA’s essential cyber security requirements has been evaluated against the legal requirements imposed by the GDPR, that is, the General Data Protection Regulation [8]. In addition, the CRA’s reporting obligations with respect to severe cyber security incidents have been analyzed in conjunction with the EU’s NIS2 directive for critical infrastructure protection [9]. There is also existing work on the CRA’s other reporting obligations with respect to vulnerability coordination and disclosure, including the mandatory reporting of actively exploited vulnerabilities [10]. The present paper continues this comparative evaluation work.
The alignment evaluation raises also the paper’s practical relevance because the regulation will be supported by standards. To this end, also European standardization and other related organizations have already conducted CRA-specific evaluation studies [11, 12]. Related work is being done by open source software communities [13] who too perceive standards and standardized processes as important for reaching compliance with the CRA [14]. While recognizing that the terms validation and evaluation have specific meanings in requirements engineering [15], the evaluation term is used because the CRA’s requirements and the ATT&CK®’s mitigations are not validated with an existing product or an implementation. It should be also mentioned that the CRA’s essential cyber security requirements are legal requirements imposed by a law. Therefore, the paper can be framed also toward existing requirements engineering research dealing with legal requirements specifically [8, 16, 17].
A further important point is that the ATT&CK®’s mitigations are categorized into three groups: enterprise, mobile, and industrial control systems. Thus, not all of the mitigations are strictly about products, which is what the CRA is mainly about, but their generality still allows to evaluate the alignment. In fact, many of the natural language descriptions for the mitigations resemble actual requirements to some extent. They are also known as countermeasures [18] or security controls [19]. It should be further emphasized that the CRA’s essential cyber security requirements are risk-based; they should be prioritized according to results from risk analyses. Even though the ATT&CK® framework could be used to help at a risk analysis, as also demonstrated in the literature [20], the paper only considers the alignment evaluation. Even under this restriction, the paper helps at identifying relevant mitigations that can support compliance with the CRA’s essential requirements at a conceptual level, thus also bridging the gap between legal and technical domains
The paper’s remainder is structured into three sections. The opening Section II elaborates the analytical evaluation methodology, presenting also research questions to guide the evaluation. Then, the evaluation results are presented in Section III. The final Section IV presents a concluding discussion.
II Approach
A terminological clarification is required to elaborate the analytical evaluation methodology and the research questions (RQs). The alignment is evaluated by using a generic concept of a mapping. Different mappings have also frequently been used to compare, evaluate, and develop standards and frameworks [5, 19, 21, 22]. Specifically, “a mapping is a bi-directional connection between information contained in two artifacts”, whereas “a link is a uni-directional connection from one artifact to another” [23, p. 67]. With this simple terminology, which closely resembles definitions used in software traceability research [24, 25], the artifacts examined refer to the ATT&CK®’s mitigations and the CRA’s essential requirements categorized into requirement groups in previous work [8]. A bidirectional connection between the two means that in case an essential cyber security requirement is picked from the CRA, a corresponding mitigation is present in the ATT&CK® framework—and the other way around. To simplify the analysis, a restriction is placed: each mitigation can only map to one essential cyber security requirement. Thus, in terms of Fig. 1, only one-to-one mappings are considered. In practice, a mapping of a mitigation was done by considering the most representative and illuminating requirement.
Figure 1: Alignment, Gaps, and Mappings
Because the CRA is a law and the MITRE’s ATT&CK® is a framework, a prior hypothesis is that at least some of the essential requirements map to multiple mitigations because legal requirements cannot be specified in detailed, technical terms. This point also correlates with research on the real or perceived ambiguity of many legal requirements [17, 26]. Having said that, also the ATT&CK® operates at a rather high abstraction level, meaning that fine-grained technical details are absent from the mitigations offered and suggested by the framework. In other words, standards too are often ambiguous and incoherent. For instance, there is a lack of consistency even regarding the definitions for fundamental concepts, including for the confidentiality, integrity, and availability (CIA) triad [19]. As the alignment evaluation is based on manually done mappings, which have been the de facto approach in existing work [21, 22], the ambiguities and inconsistencies also raise a validity threat, as is typical in qualitative research.
To address the threat, a collaboration process used in previous work [27] was adopted. The collaboration involved three steps: (1) the first author made the initial mappings and identified the initial gaps; (2) the other two authors reviewed these and raised arguments in case of disagreements; (3) all three authors resolved the disagreements by short negotiations. As usual, a perfect consensus holds for the final mappings and gaps—after all, otherwise an author would not be an author.
Another point to draw from Fig. 1 is that some of the requirements may not map to any mitigations, or the other way around, which would indicate a presence of a gap or several gaps. Explicit identification of gaps in elicited requirements is generally important [28], including with respect to changing requirements [29], which can be seen to frame also the CRA’s essential cyber security requirements [8]. With the previous points about legal requirements in mind, a prior expectation again is that there are some gaps in the CRA. On one hand: not everything can be covered in a law—and arguably not everything should even be covered in a law. On the other hand: if there are numerous gaps in the CRA vis-à-vis the ATT&CK®, it could be argued that the policy-makers overlooked some important aspects—a criticism that has been expressed in relation to funding or other support for open source software projects and communities [27]. In any case, the overall alignment can be summarized by counting the gaps, one-to-one mappings, and one-to-many mappings.
With these elaborations, the following RQs are evaluated:
  • RQ.1: How well the CRA aligns with the ATT&CK®?
  • RQ.2: How many and what kinds of gaps there are?
Before continuing to the results, a couple of additional remarks are warranted about the manually but collaboratively done mappings. The first is that the CRA’s essential cyber security requirements were mapped to the mitigations by using the twelve collated non-functional requirement groupings elicited and conceptualized in previous work [8, 16]. These simplify the presentation because no explicit legal references are required. However, the second point follows: the CRA’s essential requirements contain also the CIA triad, which can be argued to be related, either explicitly or implicitly, to all other essential requirements. Initial disagreements were present particularly with respect to the CIA triad’s relation to the CRA’s essential cyber security requirement to review and apply exploitation mitigation techniques. As said, these were resolved through negotiations between the authors. A guiding principle for the negotiations was about the earlier remark about mapping the mitigations to the “most representative and illuminating requirement”. To apply this guideline in practice, all authors were instructed to first consider other requirements than the CIA triad, and then, if no suitable candidate was present, to use the triad for a given ATT&CK®’s mitigation.
Figure 2: The Mappings Between the CRA’s Essential Cyber Security Requirements and the ATT&CK®’s Mitigations for Enterprises
Figure 3: The Mappings Between the CRA’s Essential Cyber Security Requirements and the ATT&CK®’s Mitigations for Industrial Control Systems (rectangles colored in white denote those mitigations that are absent in Fig. 2)
III Results
The manually and collaboratively done mappings are shown in Figs. 2 and 3 for the ATT&CK®’s mitigations for enterprises and industrial control systems, respectively. Before continuing to unpack these mappings, a remark should be made: as the ATT&CK® framework has grown substantially throughout the years [3], the mappings apply only to the situation at the of writing, May 2025. With this point in mind, it can be started by remarking that agreement between the three authors was very good: in total, only four disagreements were recorded. In addition, five disagreements were raised by the two reviewing authors about within-CRA and within-ATT&CK® mappings, which indicates that the exclusions in Fig. 1 are too restrictive. As has already been pointed out [8], even the CRA’s requirements are related to each other.
The main conclusion to draw from the two figures is that the alignment is generally very good. Regarding the $12$ collated non-functional requirements of the CRA and the $44$ mitigations for enterprises, only three mitigations could not be mapped. These are the ATT&CK®’s mitigations to have a threat intelligence platform, to build an out-of-band communications channel for secure exchanges during incident management, as also recommended by other frameworks [30], and to provide training for users on matters such as phishing and social engineering in general. All three gaps can be seen to be on the organizational side. Also the CRA’s essential requirements about security testing and vulnerability coordination have been seen as being on an organizational side instead of the technical product-side [8]. Thus, in addition to the disagreements already noted, some ambivalence is present regarding functional and non-functional but technical requirements and organizational cyber security requirements.
Regarding the ATT&CK®’s mitigations for industrial control systems, the gaps are the same but with an addition of mitigations that cannot be reasonably implemented. This “mitigation limited or not effective” category connotes with a concept of residual risk in cyber security risk management research and practice [30, 31]. In other words, a risk—as a probabilistic concept, still remains after a design and an implementation of mitigations. Otherwise, the mitigations in Fig. 3 are more comprehensive than the mitigations for enterprises. The CIA triad stands out in this regard. Unlike with the enterprise mitigations, also supply chain security is accounted for. It maps to the CRA’s obligation to establish a machine-readable software bill of materials (SBOM). SBOMs are either required or recommended also by other cyber security laws and frameworks [22, 32]. That said, the two mitigation categories are related also in a sense that the CRA is envisioned to reduce supply chain management costs, risk management costs, and incident management costs for users, including enterprises, using the products covered by the law [33]. Also a few other mitigations for industrial control systems can be explicitly mentioned. Among these are recommendations to establish safety segmentation and mechanical or physical protection layers, to minimize unnecessary wireless signal propagation, to prefer predefined allowlists regarding network destinations to which a device can connect, and to ensure availability by means such as watchdog timers and redundancy solutions.
Finally, regarding the CRA’s essential requirements, there are three visible gaps: data minimization, data erasure, and vulnerability coordination. None of these legal requirements can be seen to directly and explicitly map to any of the mitigations. This observation supports an argument raised earlier in the literature [5] about a need to continuously update the ATT&CK® framework, including with respect to mitigations and countermeasures. This updating point notwithstanding, as said, the framework aligns well with the CRA in overall.
IV Conclusion
The paper presented an alignment evaluation of the Cyber Resilience Act and the ATT&CK® framework’s mitigations for enterprises and industrial control systems. The conclusion is clear: the two align generally well with each other (RQ.1). This conclusion can be seen to also raise the CRA’s legitimacy in a sense that its essential cyber security requirements are well-known means to improve the cyber security of network-connected products. With respect to the CRA, only data minimization, data erasure, and vulnerability coordination are absent among the ATT&CK®’s mitigations considered (RQ.2). With respect to the mitigations, there are also some but not many gaps (RQ.2). Among these gaps is a recommendation related to residual risks. Since the CRA is a risk-based regulation, meaning that also its essential cyber security requirements should be prioritized according to risks identified, the residual risk concept is worth explicitly singling out from the gaps.
Regarding future research, (a) it seems sensible to further continue alignment studies with alternative cyber security frameworks [34] and particularly standards. Such studies are important because following the upcoming harmonized standards will in most cases provide a presumption of conformity and compliance with the CRA, as is typical also in many other domains in the European Union [35, 36]. However, alignment evaluations are not sufficient alone; (b) once the standards have been made, also they should be evaluated, including with respect to their practical usefulness, preciseness and clarity [12], rigor, scope, complexity, and other related evaluation criteria for standards. In a similar vein, (c) it seems that empirical evaluations are mostly absent with respect to the ATT&CK®’s usefulness for practitioners. As was noted, the framework operates at a rather high abstraction level, which may—or may not—decrease its practical usefulness.
Regarding these three paths for future research, the concepts of layering and layers could be used to move beyond mappings and alignments—a high-level concept, such as a legal cyber security requirement, should ideally descend into lower level layers, including concrete, technical layers [18]. This point also reiterates the importance of standards, which too should arguably be specific enough but still applicable to a wide range of products. Reaching a good balance in this regard may even be crucial for a success of the CRA in improving the cyber security of network-connected products in the future.